OpenStudy (anonymous):

Solve 13 – x = 4. A. –17 B. 9 C. –9 D. 17

OpenStudy (anonymous):

I THINK ITS B @UsukiDoll @dan815 @Rina.r

OpenStudy (anonymous):

@tHe_FiZiCx99 is it B???

OpenStudy (anonymous):

-9

OpenStudy (xishem):

It's B, 9. \[13-9=4\]

OpenStudy (anonymous):

dang its (C)

OpenStudy (anonymous):

si i was right B

OpenStudy (rina.r):

13-x=4 x=4=-13 x=-9 so its c

OpenStudy (anonymous):

oh its B

OpenStudy (shamil98):

13 -x = 4 13 = x+ 4 9 = x

OpenStudy (shamil98):

You did your math wrong..

OpenStudy (anonymous):

well they say C OR D LMAO WITCH ONE IS IT

OpenStudy (shamil98):

ITS B

OpenStudy (anonymous):

so wat is it

OpenStudy (rina.r):

its -9 b coz u bring 13 to the right then sign changes to -13 so subtract 4-13 gives u -9

OpenStudy (rina.r):

its C -9

OpenStudy (anonymous):

ok next question

OpenStudy (the_fizicx99):

Its B. Subtract 13-4 = 9 13-9 = 4

OpenStudy (rina.r):

bring it on

OpenStudy (anonymous):

Solve 5x = –25. A. –15 B. –4 C. –25 D. –5

OpenStudy (the_fizicx99):

Inverse. 5x/5 = -25/5 is?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

yes its D thats wat i got

OpenStudy (the_fizicx99):

@ziqbal103 not suppose to give the answer out. Your suppose to guide him.

OpenStudy (rina.r):

its D -5

OpenStudy (anonymous):

next question

OpenStudy (the_fizicx99):

Make new post. Thread gets slow.

OpenStudy (anonymous):

the- why you donn't say others and its muliple choice so i said -5

OpenStudy (usukidoll):

w t heck am I doing here?